OpenStudy (anonymous):

What is the solution of the matrix equation [9,4_2,1] X= [-9,-6_-1,-8]

OpenStudy (zarkon):

whats with the underscores

OpenStudy (zarkon):

you mean ... \[\left[\begin{array}{cc} 9 & 4 \\ 2 & 1 \\ \end{array}\right]X=\left[\begin{array}{cc} -9 & -6 \\ -1 & -8 \\ \end{array}\right]\]?